shift summary, Evening 09/18/99

Patrick Toale (toale@cuhep.Colorado.EDU)
Sat, 18 Sep 1999 23:00:12 -0600 (MDT)

Shift Summary for eve shift, 18/Sept/99

On Shift: Pat T., Nick S.

Hours of beam available at all intensities: 8.5
Hours of beam available at nominal intensity: 8.5
Hours of E799 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): 8.0 (trigger studies)

Accounting of lost time: changing runs, changing magnet settings

========================================================================
Runs ended this shift:
Run 14611 - 0.80M events - BEAM_799
Run 14612 - 0.43M events - BEAM_799
Run 14613 - 0.61M events - BEAM_799
Run 14614 - 0.76M events - BEAM_799
Run 14615 - 0.83M events - BEAM_799
Run 14616 - 0.83M events - BEAM_799
Run 14617 - 0.85M events - BEAM_799
Run 14618 - 0.89M events - BEAM_799
Run 14619 - 0.24M events - BEAM_799
Run 14620 - 0.84M events - BEAM_799
Run 14621 - 0.86M events - BEAM_799
Run 14622 - 0.83M events - BEAM_799
Run 14623 - 0.80M events - BEAM_799

Tapes brought to FCC this shift: none
(including tapes for online split)

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
E799 datataking,
beam position studies by Rick C.,
trigger rate studies at KICK=150,200 and absorber=0",10",
trigger 10 rate studies.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14611 15:30 Inherit shift.
18 15:37 Stop run.

14612 15:44 Start run.
12 16:00 Stop run.

14613 16:05 Start run.
13 16:23 Stop run.

14614 16:28 Start run.
16:40 Change magnet settings/absorber lenght to look
at rates at 150MeV/10"Be, 150MeV/0"Be,
200MeV/10"Be, and 200MeV/0"Be.
30 17:09 Stop run.

14615 17:18 Start run with KICK=150MeV and 10" Be.
32 18:02 Stop run.

14616 18:05 Start run with KICK=150MeV and 0" Be.
21 18:35 Stop run.

14617 18:41 Start run with KICK=200MeV and 0" Be.
20 19:08 Stop run.

14618 19:13 Start run with KICK=200MeV and 10" Be.
4 19:20 Problems with MF2, big PMT drifts,
cycle MF2 power, problem gone at spill 9.
31 19:54 Problem again with MF2, goes away after 2 spills,
laser events on KDISP look good.
41 20:07 Stop run.

14619 20:14 Start run with KICK=200MeV and 10" Be.,
Trig 10 now has HA in veto.
17 20:35 Stop run to change trig 10 definition.

14620 20:42 Start run with KICK=200MeV and 10" Be.,
and more changes to trig 10 (`97 definition).
42 21:40 Stop run.

14621 21:45 Start run with KICK=200MeV and 0" Be.,
and `97 trig 10 definition.
26 22:20 Stop run.

14622 22:28 Start run with KICK=150MeV and 0" Be.,
and `97 trig 10 definition.
24 22:59 Stop run.

14623 23:04 Start run with KICK=150MeV and 10" Be.,
and `97 trig 10 definition.
40 23:57 Stop run.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Onsplit debugging by Arun and gang.